OpenStudy (anonymous):

Solve for x: |x| − 8 = −5 x = −13 and x = −3 x = 3 and x = −3 x = 3 and x = 13 No solution

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

he is offline but i think i can help

OpenStudy (anonymous):

it's no solution

OpenStudy (anonymous):

Okay thank you!

OpenStudy (anonymous):

np can i get a medal and let me know if its correct

OpenStudy (anonymous):

Yeah I have one more though!

OpenStudy (anonymous):

x - 8 = -5 x - 8 + 8 = -5 + 8 x = -5 + 8 x = 3

OpenStudy (anonymous):

Hope this helped!

OpenStudy (anonymous):

Wait so what does the other x equal ?

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

it's no solution because in the answer choices only one fits in there

OpenStudy (anonymous):

No solution was wrong.

OpenStudy (anonymous):

and 13 - 8 = 3 not -5 so C and B are not correct

OpenStudy (anonymous):

and A is also not correct so it's no solution D

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

XD

OpenStudy (anonymous):

i just answered a question like this on flvs and i got it right thats why i though it was no solution

OpenStudy (anonymous):

Well it was wrong...

OpenStudy (anonymous):

are you flvs?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

which course is this?

OpenStudy (anonymous):

Im in flvs also

OpenStudy (anonymous):

Its math silly you!

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

well i have another one up if you can help me out ?

OpenStudy (anonymous):

I finished flvs about a month ago haha

OpenStudy (anonymous):

I had to restart ninth grade, a lot of crazy things have happened this year :'(

OpenStudy (anonymous):

i know it's math lol but algebra geometry or what?

OpenStudy (anonymous):

aw what happened? failing grades? :(

OpenStudy (anonymous):

No, I was actually in school but bullying started, i went to flvs for a bit, moved in with my dad and started going to a different high school, and while i was there, five suicides happened and four of them happened to be my friends.... and a lot of other stuff.

OpenStudy (anonymous):

Moral of the story, stay in home school kids, the real world hurts...

OpenStudy (anonymous):

omg i feel so bad for you how old are you?

OpenStudy (anonymous):

People these days ;( they bully because they get jealous at least you know you are better than them :)

OpenStudy (anonymous):

Im 15.

OpenStudy (anonymous):

I was bullied once its the worst but always keep in mind you are better than them.

OpenStudy (anonymous):

they usually try to put others down because they know they can't make it to the top so if the put other down they feel like they are in the top and i'm so sorry no teenager should go through this actually no one should

OpenStudy (anonymous):

Yeah Im keeping my head up though! Could you guys help me with my other question lol ?

OpenStudy (anonymous):

sure lol

OpenStudy (anonymous):

of course hopefully i get it right